In order to correct the vision of a farsighted person, a converging lens of focal length 33.8 cm is used to bring the person's near point in to the normal distance of 25 cm. Calculate the uncorrected near point of this farsighted person when eyeglasses are not used, in units of cm.
Answer #1

Solution:

Using lens formula:

\frac{1}{d_{i}}-\frac{1}{d_{o}}=\frac{1}{f}

\frac{1}{d_{i}}-\frac{1}{-25cm}=\frac{1}{+33.8cm}

d_{i}=-96.0 cm

The uncorrected near point of the farsighted person is -96.0 cm